Problem 14-03A a-d On January 1, 2020, Carla Vista Corporation had the following stockholders equity accounts. Common StockPrepare the stockholders equity section of the balance sheet at March 31. (Enter account name only and do not provide descriPrepare the stockholders equity section of the balance sheet at June 30. (Enter account name only and do not provide descripPrepare the stockholders equity section of the balance sheet at September 30. (Enter account name only and do not provide dePrepare the stockholders equity section of the balance sheet at December 31, 2020. (Enter account name only and do not provi
Answer #1

Part A

CARLA VISTA CORPORATION

Partial Balance Sheet

March 31, 2020

Stockholders' Equity

Paid-in Capital

Capital Stock

Common Stock

1375000

Retained Earning

456000

Total Stockholders' Equity

$1831000

541000-(85000*1) = 456000

Part B

CARLA VISTA CORPORATION

Partial Balance Sheet

June 30, 2020

Stockholders' Equity

Paid-in Capital

Capital Stock

Common Stock

1375000

Retained Earning

456000

Total Stockholders' Equity

$1831000

Part C

CARLA VISTA CORPORATION

Partial Balance Sheet

September 30, 2020

Stockholders' Equity

Paid-in Capital

Capital Stock

Common Stock

1528000

Retained Earning

303000

Total Stockholders' Equity

$1831000

1375000+((85000*3)*6%*10)) = 1375000+153000 = 1528000

456000-153000

Part D

CARLA VISTA CORPORATION

Partial Balance Sheet

December 31, 2020

Stockholders' Equity

Paid-in Capital

Capital Stock

Common Stock

1528000

Retained Earning

566850

Total Stockholders' Equity

$2094850

1375000+((85000*3)*6%*10)) = 1375000+153000 = 1528000

((85000*3)+(85000*3*6%)) = 270300

303000+399000-(0.50*270300)